Money is invested in a "guaranteed fund" at 12% simple interest. After one year, there is 840 in your account. How much was originally invested?
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

X*.12X=840
1.12X=840
X=840/1.12
X=$750 WAS THE ORIGINAL INVESTMENT.

Answer by stanbon(75887)

Money is invested in a "guaranteed fund" at 12% simple interest. After one year, there is 840 in your account. How much was originally invested?
----
A(t) = P(1+rt)
840 = P(1+0.12*1)
P = 840/1.12
P = $750
==============
Cheers,
Stan H.